OpenStudy (anonymous):

Assume that x and y are differentiable functions of t. A point is moving along the graph of the equation y=3x^2 -7x At what rate is y changing when x=8 and is changing at a rate of 3 units/sec? 2 hours ago - 4 days left to answer. Additional Details a. 144 units/sec b. 41 units/sec c. 51 units/sec d. 123 units/sec e. 27 units/sec

OpenStudy (lgbasallote):

-4 days left to answer? that's new...

OpenStudy (kainui):

So all you're doing here is differentiating with respect to t instead of x as normal. Think as though you're doing implicit differentiation! I'll go ahead and show you: \[\frac{ d }{ dt } (y)=\frac{ d }{ dt } (3x^2-7x)\] \[\frac{ dy }{ dt } =6x\frac{ dx }{ dt }-7\frac{ dx }{ dt } \] Now you're given dx/dt and x, just solve for dy/dt like the problem asks for algebraically! Problem solved.